The directions are Use the Pythagorean Theorem to find the missing side. Then calculate the perimeter.
We have a trapezoid. Top base is 30 and bottom base is 58. The height is 50. Need to find the two legs which are each X. There are no angles mentioned.
Thanking you so much in advance.

Answer by checkley77(12844)

THE ENDS ARE THE HYPOTENUSE OF A 14,50,X RIGHT TRIANGLE.
14^2+50^2=X^2
196+2,500=X^2
X^2=2,696
X=SQRT2,696
X=51.923 ANS FOR THE END SIDE.
P=58+30+51.923+51.923
P=191.846 ANS. FOR THE PERIMETER.